Customers arrive at a shopping mart randomly at a rate of 7 per hour.

a)     Find the probability that during any 90 minutes’ period, the number of customers arriving at the shopping mart is exactly 8.

b)     If a customer arrives at 11.30 am, find the probability that the next patient arrives before 11.45 am.


1
Expert's answer
2022-04-05T15:54:35-0400

a

"\lambda"=7 per hour

in 90 mins "\lambda"=10.5 per hour

P(x)="\frac{\lambda^x(e^-\lambda)}{x!}"

P(8)="\frac{10.5^8\cdot e^-10.5}{8!}"

P(8)=0.1009

b

For 15 mins "\lambda"=7/4 per 15 mins

P(1)="\frac{(7/4)^1\cdot e^-7/4}{1!}"

P(1)=0.3041